# Hartman Effect from a Geometrodynamic Extension of Bohmian Mechanics

### v2

@@ Abstract
This paper presents the derivation of a general solution to the scattering problem of particles incident onto a barrier of constant potential. This solution is constructed through a geometrodynamic approach to Bohmian mechanics, assuming that particles undergo quantum tunneling along geodesic trajectories in an Alcubierre-like spacetime. Furthermore, from this solution, mathematical expressions for the quantum potential, momentum, position, and tunneling time are determined in terms of the spacetime geometry for each relevant region. This allows us to explain the Hartman effect as a consequence of spacetime distortion generated by the quantum potential within the barrier.
@@ 1 Introduction
The unusual saturation behavior in tunneling times discovered by Hartman in 1962, known as the Hartman Effect [1], exhibited by particles incident upon a very wide potential barrier [2, 3, 4], has led to discussions about the possibility of superluminal tunneling speeds [5, 6, 7, 8]. Initially, this sparked a vigorous debate regarding whether these results were due to inaccuracies in tunneling time measurements [9]. Subsequently, it compelled numerous theorists of the time to embark on the intricate task of addressing the fundamental question: What is the correct definition of tunneling time [10, 11, 12, 13, 14, 15, 16, 17]?
@@
This question, captivating not only for physicists of that era but also relevant to contemporary researchers, has endured through the years, with many attempts to answer it proving unsuccessful across various approaches [13, 14, 18, 19]. However, despite unsuccessful attempts, the Hartman effect has led to interesting applications that propose taking advantage of this effect in sending signals (information) faster than in the case of free evolution, which is impossible due to an apparent trade-off between speedup and intensity caused by the exponential damping due to the tunnel [20].
@@
Despite the considerable body of literature addressing this age-old paradox, existing work primarily focuses on definitions derived from approaches in non-relativistic and relativistic quantum mechanics. As previously mentioned, these approaches fall short of providing a satisfactory explanation for the observed effect [21, 22, 23, 11, 24, 25, 26, 12, 13, 18, 14, 27, 19, 28]. In contrast to these previous approaches, within the context of Bohmian mechanics, the question of defining tunneling time has not been sufficiently addressed [29, 30], despite a considerable body of literature on the application of Bohmian theory to the particle scattering problem over a potential barrier [30]. One of the more intriguing works in this context addresses a similar question and investigates Klein tunneling times for electrons in a two-terminal graphene device comprising a potential barrier between two metallic contacts [14, 31].
@@
On the other hand, due to the mathematical structure of Bohmian mechanics [32, 33, 34], it is possible, through certain physical and mathematical considerations, to extend it to a version that incorporates elements of spacetime as defined in general relativity [35, 36]. Therefore, it becomes feasible to develop a geometrodynamic description of quantum systems and to derive mathematical expressions that are not provided by approaches based on the orthodox formalism of relativistic or non-relativistic quantum theory [37]. Consequently, this work presents the construction of a solution that explains the Hartman Effect through a geometrodynamic approach to Bohmian mechanics, considering a spacetime endowed with a metric that allows for superluminal travel without violating the principles of general relativity [38].
@@
More specifically, the present work reports a general solution constructed under the hypothesis that particles, during the tunneling phenomenon, follow geodesic trajectories within an Alcubierre-like spacetime [38, 39]. Consequently, from this solution, quantities of interest are determined, including mathematical expressions describing linear momenta, quantum potentials, and particle positions for the problem under investigation. As a result, the particle’s position expression leads to the mathematical expression of the tunneling time, equipped with the geometric elements of the considered spacetime.
@@
The remainder of this article is organized as follows. Section 2 provides a brief yet necessary review of the fundamental concepts of Bohmian mechanics. Section 3 deduces the geometrodynamic version of the quantum force based on the results presented in the literature[37]. Section 4 presents the construction of the general solution to the problem of a particle incident from the left on a barrier of constant potential for each of the regions of interest. In Section 5, the general expression of the quantum potential for each region of interest is determined. Section 6 establishes the general expression for linear momentum and particle position, again for each region of interest, consequently providing the mathematical expression for the tunneling time. Section 7 presents the conclusion of the work. We also include an Appendix containing some mathematical results and methods used in the main text.
@@ 2 Fundamentals of Bohmian mechanics
Bohmian mechanics is a theory discovered by Louis de Broglie in 1927 and later rediscovered by David Bohm in 1952. It provides a description of quantum systems through a mathematical object known as the wave function, which encodes their dynamics [33]. Essentially, the partial description of the quantum system is given by a wave function that evolves according to the Schrödinger equation.
@@
where $\rho\equiv\rho(\vec{x},t)$ and $S\equiv S(\vec{x},t)$. Upon substituting this function into the Schrödinger equation,
@@
$-\frac{\hbar^{2}}{2m}\frac{\nabla^{2}\sqrt{\rho}}{\sqrt{\rho}}+\frac{\hbar^{2}}{2m}(\nabla S)^{2}+V=-\hbar\partial_{t}S$ (3)
@@
$-\frac{\hbar^{2}}{2m}\biggl{[}2(\nabla\sqrt{\rho})\cdot\nabla S+\sqrt{\rho}(\nabla^{2}S)\biggr{]}=\hbar\partial_{t}\sqrt{\rho}.$ (4)
@@
$m\partial_{t}\vec{v}+m(\vec{v}\cdot\nabla)\vec{v}=-\nabla(V+Q)=\vec{F}+\vec{F}_{Q}.$ (11)
@@ 3 Geometrodynamics and Quantum Force
where $\Gamma_{\alpha\beta}^{\mu}$ are the Christoffel symbols, the indices $\mu$, $\alpha$, and $\beta$ take values in $\{0,1,2,3\}$. Equation (12) yields the geometrodynamic constraint:
@@
$\begin{split}\ddot{x}^{i}=d_{t}^{2}x^{i}+G^{i}_{jk}\partial x^{j}\partial x^{k}=d_{t}^{2}x^{i}+\omega^{i}.\end{split}$ (18)
Above $G_{jk}^{i}$ are the Christoffel symbols in the orthogonal coordinate system. Then, considering the expressions (16), (17), and (18), equation (12) is reformulated as:
$\begin{split}\Gamma^{i}_{jk}d_{t}x^{j}&d_{t}x^{k}-2c\Gamma^{i}_{0j}d_{t}x^{j}+c^{2}\Gamma^{i}_{00}=\omega^{i}+\frac{g^{ij}}{m\sqrt{|g|}}\partial_{j}V+\frac{g^{ij}}{m\sqrt{|g|}}\partial_{j}Q.\end{split}$ (19)
Next, considering the general expression for the momentum components, $(\vec{P})^{i}=\hbar(\nabla S)^{i}$, which, similarly to (16) or (17), can be rewritten as
@@
$\begin{split}\Gamma^{i}_{jk}=\frac{g^{ih}}{2}\{\partial_{j}g_{hk}+\partial_{k}g_{jh}-\partial_{h}g_{jk}\},\;\;\Gamma^{i}_{j0}=\frac{g^{ih}}{2}\{\partial_{j}g_{h0}+\partial_{0}g_{jh}-\partial_{h}g_{j0}\}\;\;\mbox{ and}\;\;\Gamma^{i}_{00}=\frac{g^{ih}}{2}\{2\partial_{0}g_{h0}-\partial_{h}g_{00}\},\end{split}$ (21)
@@
$\begin{split}&\frac{\hbar^{2}g^{ih}}{2(m\sqrt{|g|})^{2}}\{\partial_{j}g_{hk}+\partial_{k}g_{jh}-\partial_{h}g_{jk}\}g^{jl}g^{kn}\partial_{l}S\partial_{n}S-\frac{c\hbar g^{ih}}{m\sqrt{|g|}}\{\partial_{j}g_{h0}+\partial_{0}g_{jh}-\partial_{h}g_{j0}\}g^{jl}\partial_{l}S\\ &\hskip 192.05609pt+\frac{c^{2}g^{ih}}{2}\{2\partial_{0}g_{h0}-\partial_{h}g_{00}\}=\omega^{i}+\frac{g^{il}}{m\sqrt{|g|}}\partial_{l}V+\frac{g^{il}}{m\sqrt{|g|}}\partial_{l}Q.\;\end{split}$ (22)
@@ 4 Construction of the General Solution to the Problem
In this section, we present the deduction of the Bohmian solution to the one-dimensional problem of the scattering of non-relativistic and massive particles incident from the left over a barrier of constant potential [42, 3, 2], based on the direct application of the formalism introduced in the previous sections 2 and 3. Specifically, what is presented is the construction of an ansatz composed of the solution for each of the regions of interest, offered by the formalism’s of the sections 2 and 3, and related between them by the conditions of continuity and under the protection of the local equivalence relationship between the Euclidean space-time, where Bohm mechanics is defined, and the space-time of the Lorentz variety offered by Eq. (13).
@@
$\Theta(x)=\begin{cases}0&\mbox{for }x<0\\ 1&\mbox{for }x\geq 0\end{cases}.$ (24)
@@
Figure 1: Schematic representation of the quantum tunneling phenomenon of a particle with energy $E$, incident from the left on a potential barrier of height $V_{0}$ and width $a$. Specifically, images (a), (b) and (c) are presented, where the regions of interest $I$, $II$ and $III$ are considered. In image (a), the incidence of a Bohmian particle in black and its guide wave in yellow on the potential barrier is presented in region $I$, in image (b), in region $II$ the Bohmian particle in black and its guiding wave in blue, inscribed within the hypothetical Alcubierre ball, a larger circumference with a discontinuous line, that is, the hypothetical representation of the tunnel effect as a distortion of space-time. Finally, in (c) the transmitted Bohmian particle in black and its guide wave in red. In that sense, the regions $I$, $II$ and $III$ are considered in the construction of the general solution presented in section 4.
So, the ansatz that describes the motion of particles in regions $I$ and $II$ must be of the form $\psi\propto\exp{\{-ip_{\mu}x^{\mu}/\hbar\}}=\exp{\{-i[kx-Et]/\hbar\}}=\exp{\{iS(x,t)\}}$, which satisfies the Klein-Gordon equation
@@
The dispersion relation[46] is $E=\pm c\sqrt{m_{0}^{2}c^{2}+k^{2}}$. In the non-relativistic limit, where $E^{\prime}=E-m_{0}c^{2}$, the wave function takes the form $\psi\propto\phi(x,t)\exp{\{-im_{0}c^{2}t/\hbar\}}$. Thais leads to the Schrödinger equation
@@
$\begin{split}\psi_{I}(x,t)&=A\exp{\{ip_{\mu}x^{\mu}/\hbar\}}+B\exp{\{-ip_{\mu}x^{\mu}/\hbar\}}=\sqrt{\rho_{I}}\exp{\{iS_{I}(x,t)\}}\ \text{for}\ x\leq-a/2\end{split}$ (27)
and
$\begin{split}\psi_{III}(x,t)&=F\exp{\{ip_{\mu}x^{\mu}/\hbar\}}=\sqrt{\rho_{III}}\exp{\{iS_{III}(x,t)\}}\ \text{for}\ x\geq a/2,\end{split}$ (28)
where $\sqrt{\rho_{I}}=\sqrt{\mathcal{A}^{2}\cos^{2}{\{p_{\mu}x^{\mu}/\hbar\}}+\mathcal{B}^{2}\sin^{2}{\{p_{\mu}x^{\mu}/\hbar\}}}$ with $\mathcal{A}=A+B$ and $\mathcal{B}=i(A-B)$, $\sqrt{\rho_{III}}=F=\mathcal{F}$, $S_{I}(x,t)=\tan^{-1}{\{\mathcal{B}\tan{\{p_{\mu}x^{\mu}/\hbar\}}/\mathcal{A}\}}$, and $S_{III}(x,t)=p_{\mu}x^{\mu}/\hbar$. However, to obtain a consistent Bohmian solution for this problem, as shown elsewhere[33], we can write the general solution for regions $I$ and $III$ as a superposition of solutions with different energy eigenstates:
$\begin{split}\Psi_{j}(x,t)&=\sum_{l=1}^{2}C_{j}^{l}\psi_{j}^{l}(x,t)=\sum_{l=1}^{2}C_{j}^{l}\sqrt{\rho_{j}^{l}}\exp{\{iS_{j}^{l}(x,t)\}},\end{split}$ (29)
where $j=I$ for $x\leq-a/2$ and $j=III$ for $x\geq a/2$. Next, the general solution can be rewritten in the form of the Bohmian wave function, where the phase functions depending on position and time are given by:
@@
$\begin{split}&\frac{c^{2}g^{10}\partial_{0}g_{00}}{2}+\frac{c^{2}g^{11}\{2\partial_{0}g_{10}-\partial_{1}g_{00}\}}{2}=-\frac{g^{11}\partial_{1}Q}{m\sqrt{|g|}}\;\;\mbox{for}\;i=1,\ j\neq 1,\\ &\frac{c\hbar g^{10}(\partial_{1}g_{00})\partial_{1}S}{m\sqrt{|g|}}-\frac{g^{11}\partial_{1}Q}{m\sqrt{|g|}}=-\frac{g^{11}\partial_{1}Q}{m\sqrt{|g|}}\;\;\mbox{for}\;i=j=1,\ k\neq 1,\\ &\frac{\hbar^{2}g^{01}(2\partial_{1}g_{01})(g^{11})^{2}(\partial_{1}S)^{2}}{2(m\sqrt{|g|})^{2}}-\frac{c\hbar g^{10}(\partial_{1}g_{00})g^{11}\partial_{1}S}{m\sqrt{|g|}}-\frac{g^{11}\partial_{1}Q}{m\sqrt{|g|}}=-\frac{g^{11}\partial_{1}Q}{m\sqrt{|g|}}\;\;\mbox{for}\;i=j=k=1.\;\end{split}$ (32)
@@
$g_{\mu\nu}$ =($=\begin{pmatrix}-\left\{\frac{3Q}{mc^{2}}\right\}^{\frac{2}{3}}&\mathcal{O}_{1x3}\\ \mathcal{O}_{3x1}&\mathcal{I}_{3x3}\end{pmatrix},$ (33)
@@
$g^{\mu\nu}$ =($=\begin{pmatrix}-\left\{\frac{3Q}{mc^{2}}\right\}^{-\frac{2}{3}}&\mathcal{O}_{1x3}\\ \mathcal{O}_{3x1}&\mathcal{I}_{3x3}\end{pmatrix},$ (34)
where $\mathcal{I}_{3x3}$ is the identity matrix of order $33x$ 3, $\mathcal{O}_{3x1}$ the null column matrix of order $13x$ 1 and $\mathcal{O}_{1x3}$ is the null line matrix elements of order $31x$ 3.
On the other hand, in region $II$ the metric that describes spacetime must be of the Alcubierre type [38], of the Krasnikov type [47], or some type of metric that describes a hyperbolically separable spacetime, since these types of metrics support particle travel at hyperluminous speeds [48]. In this sense, in the present work the Alcubierre-like metric given by
$\mathcal{G}_{\mu\nu}$ =(ß$=\begin{pmatrix}\mbox{\ss}_{k}\mbox{\ss}^{k}-1&-\mbox{\ss}_{i}\\ -\mbox{\ss}_{j}&\gamma_{ij}\end{pmatrix},$ (35)
which has the form of the metric reported in the literature[43], where this behavior was reported and whose inverse tensor is given by the expression
$\mathcal{G}^{\mu\nu}=\begin{pmatrix}-1&-\mbox{\ss}^{i}\\ -\mbox{\ss}^{j}&\gamma^{ij}-\mbox{\ss}^{i}\mbox{\ss}^{j}\end{pmatrix},$ (36)
where $i,j,k=1,2,3$, ß$\mbox{\ss}_{i}=v_{s}(t)f(r_{s})$ for $i=1$, ß$\mbox{\ss}_{i}=0$ for $i\neq 1$, ß$\mbox{\ss}_{i}=\gamma_{ij}\mbox{\ss}^{j}$, $\gamma_{ij}=\mbox{diag}(1,1,1)$, $v_{s}\equiv v_{s}(t)=\dot{x}_{s}(t)$, $f(r_{s})=\alpha_{0}/\{v_{s}[1+\alpha_{1}\cosh{(2\sigma r_{s})}]\}$, $r_{s}\equiv r_{s}(t)=x-x_{s}(t)$, and $x_{s}(t)=nct=nt$.

### v3

@@ Abstract
This paper develops a geometrodynamic extension of Bohmian mechanics to describe quantum tunneling through a potential barrier, treating particle trajectories as geodesics in an Alcubierre-type spacetime. The model provides analytical expressions for the quantum potential, particle dynamics, and tunneling time, explicitly linked to the underlying spacetime geometry. For narrow barriers, the tunneling time depends on the barrier width, while for sufficiently wide barriers, it saturates to a constant value—recovering the Hartman effect. This behavior arises from a geometric self-regulation mechanism, where the quantum potential dynamically adjusts the spacetime distortion to maintain a fixed tunneling time, consistent with relativistic causality despite effective superluminal propagation. The results establish a direct connection between quantum tunneling and spacetime geometry, offering a unified framework to interpret the Hartman effect. This approach naturally incorporates relativistic constraints while suggesting that similar geometric mechanisms may underlie other quantum phenomena, such as topological phases in condensed matter systems.
@@ 1 Introduction
The unusual mathematical expression of a tunneling time independent of the barrier width, discovered in Ref. [1] and exhibited by particles incident on a very wide potential barrier [2, 3, 4], led to discussions about the possibility of superluminal tunneling velocities [5, 6, 7, 8, 9]. Initially, this triggered a vigorous debate about whether this result, later termed the Hartman effect, was due to inaccuracies in the tunneling time measurements [10, 11]. It subsequently forced numerous theorists to address the fundamental question [12, 13, 16, 17, 18, 14, 15, 19]: What is the correct definition of tunneling time?
Recent experimental work in Ref. [20] has reignited this debate by challenging the predictions of Bohmian mechanics regarding dwell times for particles tunneling into infinitely long barriers. Their findings, which show finite dwell times contrary to Bohmian predictions, highlight the ongoing complexity of this question. Similarly, Ref. [21] discussed how these measurements add empirical data to a discourse previously dominated by theory, yet the debate remains unresolved due to the lack of definitive predictions from standard quantum mechanics.
@@
This question, captivating not only for physicists in the past but also relevant to contemporary researchers, has endured through the years, with many attempts to answer it proving unsuccessful across various approaches [17, 18, 22, 23]. However, despite unsuccessful attempts, the Hartman effect has led to interesting applications, such as proposals to exploit this effect for sending signals (information) faster than in free evolution, though this is limited by an apparent trade-off between speedup and intensity due to exponential damping during tunneling [11].
@@
Despite the considerable body of literature addressing this paradox, existing work primarily focuses on definitions derived from non-relativistic and relativistic quantum mechanics. As mentioned above, these approaches do not provide a satisfactory explanation for the observed effect [13, 16, 17, 18, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. In contrast, within the context of Bohmian mechanics, the question of defining tunneling time has not been sufficiently addressed [33, 34], despite a considerable body of literature on the application of Bohmian theory to particle scattering over a potential barrier [34]. One of the more intriguing works in this context investigates Klein tunneling times for electrons in a two-terminal graphene device comprising a potential barrier between metallic contacts [18, 35].
@@
On the other hand, due to the mathematical structure of Bohmian mechanics [36, 37, 38, 39], it is possible, through certain physical and mathematical considerations, to extend it to a version that incorporates elements of spacetime as defined in general relativity [40, 41]. Therefore, it becomes feasible to develop a geometrodynamic description of quantum systems and to derive mathematical expressions not provided by orthodox relativistic or non-relativistic quantum theory [42, 43]. Consequently, this work presents the construction of a solution that explains the Hartman effect through a geometrodynamic approach to Bohmian mechanics, considering a spacetime endowed with a metric that allows for superluminal travel without violating the principles of general relativity [44, 45].
@@
More specifically, the present work reports a general solution constructed under the hypothesis that particles, during the tunneling phenomenon, follow geodesic trajectories within an Alcubierre-like spacetime [45, 46]. From this solution, quantities of interest are determined, including mathematical expressions describing linear momenta, quantum potentials, and particle positions for the problem under investigation. As a result, the particle’s position expression leads to the mathematical expression of the tunneling time, equipped with the geometric elements of the considered spacetime.
@@
The remainder of this article is organized as follows. Section 2 provides a brief yet necessary review of the fundamental concepts of Bohmian mechanics. Section 3 deduces the geometrodynamic version of the quantum force, starting from some results from the literature [43]. Section 4 presents the construction of the general solution to the problem of a particle incident from the left on a barrier of constant potential for each of the regions of interest. In Section 5, the general expression of the quantum potential for each region of interest is determined. Section 6 establishes the general expression for linear momentum and particle position, again for each region of interest, consequently providing the mathematical expression for the tunneling time. Section 7 presents the conclusion of the work. We also include an Appendix containing some mathematical results and methods used in the main text.
@@ 2 Fundamentals of Bohmian mechanics
Bohmian mechanics is a theory discovered by Louis de Broglie in 1927 and later rediscovered by David Bohm in 1952. It provides a description of quantum systems through a mathematical object known as the wave function, which encodes their dynamics [38]. Essentially, the partial description of the quantum system is given by a wave function that evolves according to the Schrödinger’s equation.
@@
where $\rho\equiv\rho(\vec{x},t)$ and $S\equiv S(\vec{x},t)$. Upon substituting this function into the Schrödinger’s equation,
@@
$\frac{\hbar^{2}}{2m}\frac{\nabla^{2}\sqrt{\rho}}{\sqrt{\rho}}+\frac{\hbar^{2}}{2m}(\nabla S)^{2}+V=-\hbar\partial_{t}S$ (3)
@@
$\frac{\hbar^{2}}{2m}\biggl[2(\nabla\sqrt{\rho})\cdot\nabla S+\sqrt{\rho}(\nabla^{2}S)\biggr]=\hbar\partial_{t}\sqrt{\rho}.$ (4)
@@
$m\partial_{t}\vec{v}+m(\vec{v}\cdot\nabla)\vec{v}=-\nabla(V+Q)=\vec{F}_{V}+\vec{F}_{Q}.$ (11)
@@ 3 Geometrodynamics and Quantum Force
where $\Gamma_{\alpha\beta}^{\mu}$ are the Christoffel symbols, the indices $\mu$, $\alpha$, and $\beta$ take values in $\{0,1,2,3\}$. The hypothesis of geodesic trajectories arises naturally when coupling quantum mechanics to general relativity, where particles follow extreme paths in a curved spacetime. This approach reconciles the principle of least action from classical mechanics with Bohmian quantum dynamics, extending it to relativistic scenarios. In this framework, equation (12) yields the geometrodynamic constraint:
@@
$\ddot{x}^{i}=d_{t}^{2}x^{i}+G^{i}_{jk}\partial x^{j}\partial x^{k}=d_{t}^{2}x^{i}+\omega^{i}.$ (18)
Above $G_{jk}^{i}$ are the Christoffel symbols in the orthogonal coordinate system. Then, considering the expressions (16), (17) and (18), and knowing that $\ddot{x}^{i}$ denotes the acceleration of the particle in the $i$-th spatial direction, calculated in curvilinear coordinates, equation (12) is reformulated as:
$\Gamma^{i}_{jk}d_{t}x^{j}d_{t}x^{k}-2c\Gamma^{i}_{0j}d_{t}x^{j}+c^{2}\Gamma^{i}_{00}=\omega^{i}+\frac{g^{ij}}{m\sqrt{|g|}}\partial_{j}V+\frac{g^{ij}}{m\sqrt{|g|}}\partial_{j}Q.$ (19)
Next, considering the general expression for the momentum components, $(\vec{P})^{i}=P^{i}=\hbar(\nabla S)^{i}$, which, similarly to Eqs. (16) and (17), can be rewritten as
@@
$\Gamma^{i}_{jk}$ $=\frac{g^{ih}}{2}\{\partial_{j}g_{hk}+\partial_{k}g_{jh}-\partial_{h}g_{jk}\},$
$\Gamma^{i}_{j0}$ $=\frac{g^{ih}}{2}\{\partial_{j}g_{h0}+\partial_{0}g_{jh}-\partial_{h}g_{j0}\},$
$\Gamma^{i}_{00}$ $=\frac{g^{ih}}{2}\{2\partial_{0}g_{h0}-\partial_{h}g_{00}\},$
@@
$\frac{\hbar^{2}g^{ih}}{2(m\sqrt{|g|})^{2}}\{\partial_{j}g_{hk}+\partial_{k}g_{jh}-\partial_{h}g_{jk}\}g^{jl}g^{kn}\partial_{l}S\partial_{n}S$
$-\frac{c\hbar g^{ih}}{m\sqrt{|g|}}\{\partial_{j}g_{h0}+\partial_{0}g_{jh}-\partial_{h}g_{j0}\}g^{jl}\partial_{l}S$ (22)
$+\frac{c^{2}g^{ih}}{2}\{2\partial_{0}g_{h0}-\partial_{h}g_{00}\}=\omega^{i}+\frac{g^{il}}{m\sqrt{|g|}}\partial_{l}V+\frac{g^{il}}{m\sqrt{|g|}}\partial_{l}Q.$
@@ 4 Construction of the General Solution to the Problem
In this section, we present the deduction of the Bohmian solution to the one-dimensional problem of the scattering of non-relativistic and massive particles incident from the left over a barrier of constant potential [2, 3, 48], based on the direct application of the formalism introduced in the previous sections 2 and 3. Specifically, what is presented is the construction of an ansatz composed of the solution for each of the regions of interest, offered by the formalism of the sections 2 and 3, and related between them by the conditions of continuity and under the protection of the local equivalence relationship between the Euclidean space-time, where Bohm mechanics is defined, and the space-time of the Lorentz variety offered by Eq. (13).
@@
$\Theta(x)=\left\{\begin{array}[]{cc}0&\mbox{for $x<0$}\\ 1&\mbox{for $x\geq 0$}\end{array}.\right$.
@@
Figure 1: Schematic representation of the quantum tunneling phenomenon of a particle with energy $E$, incident from the left on a potential barrier of height $V_{0}$ and width $a$. Specifically, images (a), (b) and (c) are presented, where the regions of interest II, IIII and $III$ are considered. In image (a), the incidence of a Bohmian particle in black and its guide wave in yellow on the potential barrier is presented in region $I$. In image (b), in region $II$ the Bohmian particle in black and its guiding wave in blue, inscribed within the hypothetical Alcubierre ball, a larger circumference with a discontinuous line, that is, the hypothetical representation of the tunnel effect as a distortion of space-time. Finally. In image (c) the transmitted Bohmian particle in black and its guide wave in red. In that sense, the regions II, IIII and $III$ are considered in the construction of the general solution presented in section 4.
So, the ansatz that describes the motion of particles in regions $I$ and $III$ must be of the form $\psi\propto\exp{\{-ip_{\mu}x^{\mu}/\hbar\}}=\exp{\{-i[kx-Et]/\hbar\}}=\exp{\{iS(x,t)\}}$, which satisfies the Klein-Gordon equation
@@
The dispersion relation [42] is $E=\pm c\sqrt{m_{0}^{2}c^{2}+k^{2}}$.
In the non-relativistic limit, where $E^{\prime}=E-m_{0}c^{2}$, the wave function takes the form $\psi\propto\phi(x,t)\exp{\{-im_{0}c^{2}t/\hbar\}}$. This leads to the Schrödinger equation
@@
$\psi_{I}(x,t)$ $=$ $A\exp{\{ip_{\mu}x^{\mu}/\hbar\}}+B\exp{\{-ip_{\mu}x^{\mu}/\hbar\}}$
$=$ $\sqrt{\rho_{I}}\exp{\{iS_{I}(x,t)\}}\$
for $x\leq-a/2$ and
$\psi_{III}(x,t)$ $=$ $F\exp{\{ip_{\mu}x^{\mu}/\hbar\}}$
$=$ $\sqrt{\rho_{III}}\exp{\{iS_{III}(x,t)\}}\,$
for $x\geq a/2$, where $\sqrt{\rho_{I}}=\sqrt{\mathcal{A}^{2}\cos^{2}{\{p_{\mu}x^{\mu}/\hbar\}}+\mathcal{B}^{2}\sin^{2}{\{p_{\mu}x^{\mu}/\hbar\}}}$ with $\mathcal{A}=A+B$ and $\mathcal{B}=i(A-B)$, $\sqrt{\rho_{III}}=F=\mathcal{F}$, $S_{I}(x,t)=\tan^{-1}{\{\mathcal{B}\tan{\{p_{\mu}x^{\mu}/\hbar\}}/\mathcal{A}\}}$, and $S_{III}(x,t)=p_{\mu}x^{\mu}/\hbar$. To reconcile the experimental findings of Sharoglazova et al. [20]—where finite dwell times were observed for particles tunneling into a barrier, contradicting the Bohmian prediction of infinite dwell times—the general solution in regions $I$ and $III$ must be expressed as a superposition of distinct energy eigenstates, as argued in [38, 36]:
$\Psi_{j}(x,t)=\sum_{l=1}^{2}C_{j}^{l}\psi_{j}^{l}(x,t)=\sum_{l=1}^{2}C_{j}^{l}\sqrt{\rho_{j}^{l}}\exp{\{iS_{j}^{l}(x,t)\}},$ (28)
where $j=I$ for $x\leq-a/2$ and $j=III$ for $x\geq a/2$. Here, $S_{j}^{1}(x,t)$ and $S_{j}^{2}(x,t)$ correspond to phase functions associated with two distinct energy eigenvalues $E^{1}$ and $E^{2}$, respectively, introduced to model the interference and finite dwell time observed experimentally [20]. This superposition is necessary to describe:
  • 1.
    The interference between incident and reflected waves (Region I),
  • 2.
    Nontrivial transmission (Region III),
  • 3.
    And crucially, to ensure consistency with both probability conservation and the observed energy-speed relationship $v=\sqrt{2|\Delta|/m}$ from [20], where faster motion corresponds to more negative local kinetic energy.
Rewriting the solution in the Bohmian form reveals how finite velocities emerge:
@@
The phase gradients in (29) break the Bohmian stasis paradox—where single eigenstates predict stationary particles in the barrier—by introducing non-zero velocity fields via $\nabla S_{j}(x,t)$. This resolves the discrepancy with experimental results while preserving the causal structure of Bohmian mechanics.
@@
$0$ $=$ $\frac{c^{2}g^{10}\partial_{0}g_{00}}{2}+\frac{c^{2}g^{11}\{2\partial_{0}g_{10}-\partial_{1}g_{00}\}}{2}\frac{g^{11}\partial_{1}Q}{m\sqrt{|g|}}\;\;\{\mbox{for}\;i=1,\ j\neq 1\},$
$0$ $=$ $\frac{c\hbar g^{10}(\partial_{1}g_{00})\partial_{1}S}{m\sqrt{|g|}}\;\;\{\mbox{for}\;i=j=1,\ k\neq 1\},\;$ (31)
$0$ $=$ $\frac{\hbar^{2}g^{01}(2\partial_{1}g_{01})(g^{11})^{2}(\partial_{1}S)^{2}}{2(m\sqrt{|g|})^{2}}\frac{c\hbar g^{10}(\partial_{1}g_{00})g^{11}\partial_{1}S}{m\sqrt{|g|}}\;\;\{\mbox{for}\;i=j=k=1\}.$
@@
$g_{\mu\nu}=\left(\begin{array}[]{cc}-\left\{\frac{3Q}{mc^{2}}\right\}^{\frac{2}{3}}&\mathcal{O}_{1\times 3}\\ \mathcal{O}_{3\times 1}&\mathcal{I}_{3\times 3}\end{array}\right),$ (32)
@@
$g^{\mu\nu}=\left(\begin{array}[]{cc}-\left\{\frac{3Q}{mc^{2}}\right\}^{-\frac{2}{3}}&\mathcal{O}_{1\times 3}\\ \mathcal{O}_{3\times 1}&\mathcal{I}_{3\times 3}\end{array}\right),$ (33)
where $\mathcal{I}_{3\times 3}$ is the identity matrix of order $3\times 3$, $\mathcal{O}_{3\times 1}$ the null column matrix of order $3\times 1$ and $\mathcal{O}_{1\times 3}$ is the null line matrix elements of order $1\times 3$.
On the other hand, in region $II$ the metric that describes spacetime must be of the Alcubierre type [45], of the Krasnikov type [51], or some type of metric that describes a hyperbolically separable spacetime, since these types of metrics support particle travel at hyperluminous speeds [52]. In this sense, in the present work we consider the Alcubierre-like metric, given by
$\mathcal{G}_{\mu\nu}=\left(\begin{array}[]{cc}\mbox{\ss }_{k}\mbox{\ss }^{k}-1&-\mbox{\ss }_{i}\\ -\mbox{\ss }_{j}&\gamma_{ij}\end{array}\right),$ (34)
which has the form of the metric reported in the literature [49], and whose inverse tensor is given by the expression
$\mathcal{G}^{\mu\nu}=\left(\begin{array}[]{cc}-1&-\mbox{\ss }^{i}\\ -\mbox{\ss }^{j}&\gamma^{ij}-\mbox{\ss }^{i}\mbox{\ss }^{j}\end{array}\right),$ (35)
where $i,j,k=1,2,3$ are spatial indices, $\mbox{\ss }_{i}=v_{s}(t)f(r_{s})$ for $i=1$ and $\mbox{\ss }_{i}=0$ otherwise defines the spacetime coupling along $x$, $\gamma_{ij}=\mathrm{diag}(1,1,1)$ is the flat spatial metric, $r_{s}(t)\equiv x-x_{s}(t)$ denotes the position relative to the bubble’s center, $v_{s}(t)=d_{t}[x_{s}(t)]=\dot{x}_{s}(t)$ represents the bubble’s instantaneous velocity, $f(r_{s})=\alpha_{0}/\{v_{s}[1+\alpha_{1}\cosh{(2\sigma r_{s})}]\}$ characterizes the bubble profile. And where:
  • 1.
    $\alpha_{0}=-[v_{s}\tanh{(2\sigma R)}]/[2\tanh{(\sigma R)}]$ governs the peak distortion strength;
  • 2.
    $\alpha_{1}=\mathrm{sech}{(2\sigma R)}$ controls the spatial decay rate;
  • 3.
    $\sigma$ determines the distortion region’s width;
  • 4.
    $R$ is the effective radius of the Alcubierre bubble, defining the spatial extent where the metric distortion is non-negligible ($|r_{s}|\lesssim R$).
Furthermore, the function $f(r_{s})$ peaks at $\alpha_{0}/[v_{s}(1+\alpha_{1})]$ when $r_{s}=0$ (bubble center), with $\alpha_{0}$ setting the maximum distortion amplitude. For $|r_{s}|>0$, $f(r_{s})$ decays exponentially, becoming negligible at $|r_{s}|\gg R$.